OpenStudy (anonymous):

Seven slips of paper marked with the numbers 1, 2, 3, 4, 5, 6, and 7 are placed in a box and mixed well. Two are drawn without replacement. What are the odds that the sum of the numbers on the two slips is not 6?

OpenStudy (anonymous):

is this probability?

OpenStudy (anonymous):

\[P(A) ={The~Number~Of~Ways~Event~A~Can~Occur\over The~total~number~Of~Possible~Outcomes}\]

OpenStudy (anonymous):

the event can occur 3 times, I'm just having a hard time finding all the possible outcomes

OpenStudy (anonymous):

total number of possible outcomes, i would think to be 7 since there are 7 slips of paper

OpenStudy (anonymous):

reading this could possibly help you http://www.mathgoodies.com/lessons/vol6/intro_probability.html

OpenStudy (kropot72):

There are 2 outcomes adding up to 6: 2 + 4 = 6 1 +5 = 6 P(2 and 4) =1/7 * 1/6 = 1/42 P(1 and 5) = 1/7 * 1/6 = 1/42 P((2 and 4) or P(1 and 5) = 1/42 + 1/42 = 1/21 Probability that the sum does not add up to 6 = 1 - 1/21 = 20/21 The odds are: P(sum not 6): P(sum is 6) = 20/21 : 1/21 = 20 : 1